If f(x) and g(x) are two non zero functions then prove that GCD(f(x),g(x)) = GCD(f(x)+g(x),f(x)-g(x))

Grade:12

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Let there be constant functions

f(x) = 3, g(x)= 11

gcd = 1

now f+g= 14
f-g= 8

gcd(f+g,f-g) != 1

Kindly check the questions and repost

Regards